Parameter Values for Reactor Problem

V  = 0.01  m^{3}
C_{A0}  = 1  kgmol/m^{3}
ρ  = 1000  kg/m^{3}
C_{p}  = 1  kJ/kg/K
(−ΔH_{1}) = 3.94x10^{4} kJ/kgmole
(−ΔH_{2}) = 7.48x10^{4} kJ/kgmole
(−ΔH_{3}) = 5.94x10^{4} kJ/kgmole
T_{0}  = 150  ^{o}C
k_{10}  = 100  s^{−1}
k_{20}  = 150  s^{−1}
k_{30}  = 1.1  m^{3}/(kgmole)/s
E_{1}  = 3.09x10^{4} kJ/Kgmole/K
E_{2}  = 4.18x10^{4} kJ/Kgmole/K
E_{3}  = 2.10x10^{4} kJ/Kgmole/K

• The rate constants, k_{i} are related to temperature by:

k_{i} = k_{i0}exp (−\frac{E_{i}}{RT} )

• There are 3 reactions occurring in the CSTR. The associated heats of reaction are given in the table below:

reaction heat of reaction
1 A → B (−ΔH_{1})
2 B → C (−ΔH_{2})
3 2A → D (−ΔH_{3})

• The reactions are exothermic as you can tell from the values of the heats of reaction. Thus Q will be a negative value.

Steady State Values of Inputs

\begin{aligned}F_{s} & = & 8x10^{−5} & m^{3}/s \\ Q_{s} & = & −7.5 & kJ/s \end{aligned}